OpenStudy (anonymous):

[7.04] Choose the correct simplification of the quantity of x to the 7th power over y to the 9th power all raised to the 2nd power. Select one: a. x to the 14th power over y to the 18th power b. x to the 9th power over y to the 11th power c. x to the 5th power over y to the 7th power d. x14y18

OpenStudy (anonymous):

@johnweldon1993 @BVB4rmy

OpenStudy (johnweldon1993):

\[\huge (\frac{x^7}{y^9})^2 = \frac{x^{7\times 2}}{y^{9 \times 2}} = \space ? \]

OpenStudy (anonymous):

a?

OpenStudy (johnweldon1993):

Correct again :)